1A solution of 0.235 M aspartic acid, the charge neutral form of the amino acid, is titrated with 0.118 M NaOH . The p K a values for aspartic acid are 1.990 , 3.900 , and 10.002 , corresponding to the α-carboxylic acid group, the β-carboxylic acid group, and the amino group, respectively. Calculate the pH at the first equivalence point of this titration.
Answer #1

pH in the first equivalence point,

pH = ( pKa1 + pKa2 ) / 2

In this case, where reaction is going backwards by titrating with a base,

pH = ( pKa2 + pKa3 ) / 2

= ( 3.900 + 10.002 ) / 2

= 6.95